# FedDUAL: A Dual-Strategy with Adaptive Loss and Dynamic Aggregation for Mitigating Data Heterogeneity in Federated Learning

### v1

@@ Abstract
Federated Learning (FL) marks a transformative approach to distributed model training by combining locally optimized models from various clients into a unified global model. While FL preserves data privacy by eliminating centralized storage, it encounters significant challenges such as performance degradation, slower convergence, and reduced robustness of the global model due to the heterogeneity in client data distributions. Among the various forms of data heterogeneity, label skew emerges as a particularly formidable and prevalent issue, especially in domains such as image classification. To address these challenges, we begin with comprehensive experiments to pinpoint the underlying issues in the FL training process. Based on our findings, we then introduce an innovative dual-strategy approach designed to effectively resolve these issues. First, we introduce an adaptive loss function for client-side training, meticulously crafted to preserve previously acquired knowledge while maintaining an optimal equilibrium between local optimization and global model coherence. Secondly, we develop a dynamic aggregation strategy for aggregating client models at the server. This approach adapts to each client’s unique learning patterns, effectively addressing the challenges of diverse data across the network. Our comprehensive evaluation, conducted across three diverse real-world datasets, coupled with theoretical convergence guarantees, demonstrates the superior efficacy of our method compared to several established state-of-the-art approaches. The code can be found at https://github.com/Pranabiitp/FedDUAL.
@@ 1 Introduction
Federated learning (FL) has revolutionized collaborative model training by enabling multiple clients to contribute to a global model without compromising the privacy of their local data (McMahan et al., 2017). This decentralized strategy avoids the need for sending data to a central server, thus maintaining data privacy. As the digital landscape evolves, with an increasing number of distributed data sources emerging from mobile devices, healthcare institutions, and Internet of Things (IoT) networks, FL has emerged as a pivotal solution for training sophisticated deep networks across geographically dispersed and heterogeneous environments (Bonawitz et al., 2016)Sahoo et al. (2024b)(Hu et al., 2024). However, a significant practical obstacle encountered during federated training is data heterogeneity in the form of skewness in labels and quantity of the data across various clients (Kairouz et al., 2021)(Li et al., 2020). Diverse user behaviors can lead to significant heterogeneity in the local data of different clients, leading to non-independent and identically distributed (non-IID) data. This variability can introduce biases in model training, leading to unstable convergence and potentially degrading the model’s performance or making it counterproductive (Li et al., 2022)(Zhao et al., 2018). While FedAvg (McMahan et al., 2017) is effective and widely used, it often falls short in accuracy and convergence with static aggregation methods. These methods combine model updates from different clients in a fixed manner, failing to adapt to heterogeneous data distributions and client drift, as discussed in (Karimireddy et al., 2020).
Previous studies have addressed the issue of client drift by implementing penalties for deviations between client and server models (Li et al., 2020)(Li et al., 2021a), employing variance reduction techniques during client updates (Karimireddy et al., 2020)(Acar et al., 2021), or utilizing novel aggregation methods on the server side (Chen et al., 2023)(Chowdhury & Halder, 2024).
@@ 1.1 Motivation
Prior studies by Yashwanth et al. (2024)Hu et al. (2024) have demonstrated that in non-IID scenarios, federated models tend to converge to ’sharp minima’, resulting in significant performance degradation and compromised generalizability. In this study, we investigate the root causes of this phenomenon and propose a novel solution to mitigate its effects. Our study begins with a detailed analysis of loss landscapes for FedAvg-trained models across IID and non-IID data distributions. Figure 1 visually depicts the loss landscapes of two models on the FMNIST dataset with systematic parameter perturbations. The model trained on IID data exhibits a notably smoother and wider valley in its loss landscape, suggesting greater robustness and better generalization. In contrast, the model trained on non-IID manifests sharper peaks and narrower valleys, indicating higher sensitivity to parameter variations and potential overfitting. These visualizations offer strong evidence that in the presence of non-IID data, the FedAvg algorithm achieves suboptimal generalization. Motivated by this observation, we investigate the underlying mechanisms by analyzing gradient norms to identify which parts of the neural network are most affected by data heterogeneity. Our findings, presented in Fig. 2, reveal a notable pattern: in non-IID scenarios, the gradient norms of the final layers, including the classification layer, exhibit significant amplification compared to their IID counterparts. Such amplification leads to model instability, impedes convergence, and ultimately compromises the generalizability of the federated model. Our investigation suggests that effective federated training in non-IID environments necessitates targeted adjustments during server-side aggregation, particularly for these highly affected layers, to achieve performance comparable to IID settings.
@@
This prompts one critical question: Can static aggregation methods effectively address severe non-IID data distributions across clients while maintaining higher convergence, performance, and generalizability in federated models? The answer is decidedly negative. Static aggregation methods inherently struggle with the dynamic heterogeneity present in federated networks, where adjusting parameters based on client distributions and performance in each communication round is crucial. Although incorporating predetermined parameters into the aggregation process may provide some partial mitigation, these methods fail to address the complex challenges posed by non-IID data distributions. A more dynamic and nuanced approach is necessary to effectively manage these multifaceted issues. To address this challenge, we apply dynamic aggregation to the model’s final layers, where gradient norms fluctuate significantly in non-IID scenarios, while using traditional aggregation (FedAvg) for the lower layers. For dynamic aggregation, we leverage the concept of Wasserstein Barycenter (Agueh & Carlier, 2011), derived from optimal transport theory, to integrate client-specific learning behaviors in these affected layers. By minimizing discrepancies from non-IID data, the Wasserstein Barycenter helps to align gradients from diverse clients, offering precise model updates. This approach ensures fair aggregation, adapts to data heterogeneity, reduces bias, and enhances robustness, ultimately leading to more stable model convergence and improved generalization.
@@
Figure 2: Comparison of gradient norms between models trained on IID and non-IID datasets using the FedAvg algorithm.
@@ 2 Related Work
Another line of research targets optimizing server-side aggregation in FL. For instance, Hsu et al. (2019) investigated the impact of non-IID data on visual classification by creating datasets with diverse distributions and found that increased data heterogeneity negatively affected performance, leading them to propose server momentum as a potential solution. FedNova (Wang et al., 2020b) addressed the problem of objective inconsistency due to client heterogeneity in federated optimization by introducing a normalized averaging technique, which resolves this inconsistency and ensures rapid error convergence. Addressing the limitations of traditional parameter averaging methods, Lin et al. (2020) introduced ensemble distillation for model fusion. This approach allows for the flexible aggregation of heterogeneous client models by training a central classifier on unlabeled data, using the outputs from the client models as guidance. FedMRL (Sahoo et al., 2024a) introduced a novel framework by using a loss function that promotes fairness among clients and employed a multi-agent reinforcement learning for personalized proximal terms , and a self-organizing map to dynamically adjust server-side weights during aggregation.
@@ 3 Methods and Materials
$\theta^{*}=\min_{\theta}F(\theta)=\frac{1}{K}\sum_{k=1}^{K}f_{k}(\theta)$ (1)
@@
where $p^{k}_{i}$ and $q_{i}$ are the probabilities associated with the $i^{th}$ component of the weight vectors. The local model must excel on local data while maintaining alignment with the global model to enhance overall generalization. This balance between minimizing local loss and aligning with the global model is defined as local adaptive function $\mathcal{L}_{adaptive}^{k}$ in Eq. 5.
$\mathcal{L}_{adaptive}^{k}=(1-\beta)*\mathcal{L}_{\text{local}}^{k}+\beta*D_{\text{KL}}(p^{k}\|q)$ (5)
where $\mathcal{L}_{\text{local}}^{k}$ is cross-entropy loss for $k^{th}$ client and $\beta$ is a regularization parameter and should be adaptive to account for the performance discrepancy between the local and global models. When the local model substantially outperforms the global model, $\beta$ should increase to enforce greater alignment. Conversely, if the models perform similarly, $\beta$ should decrease, allowing the local model to focus more on local optimization. The definition of $\beta$ is given in Eq. 6.
@@
$w_{k}^{t+1}=w_{k}^{t}-\eta\nabla_{w}\mathcal{L}_{\text{adaptive}}^{k}(w_{k}^{t})$ (8)
where $\eta$ is the local learning rate. Expanding the gradient term $\nabla_{w}\mathcal{L}_{\text{adaptive}}^{k}(w_{k}^{t})$, we obtain Eq. 9.
$\nabla_{w}\mathcal{L}_{\text{adaptive}}^{k}(w_{k}^{t})=(1-(\sigma(\mathcal{A}_{\text{local}}^{k}-\mathcal{A}_{\text{global}}^{k}))\nabla_{w}\mathcal{L}_{\text{local}}^{k}(w_{k}^{t})+$ (9)
$\sigma(\mathcal{A}_{\text{local}}^{k}-\mathcal{A}_{\text{global}})\nabla_{w}D_{\text{KL}}(p^{k}\|q)$
The KL divergence term, $\sigma(\mathcal{A}_{\text{local}}^{k}-\mathcal{A}_{\text{global}}^{k})\nabla_{w}D_{\text{KL}}(p^{k}\|q)$ in Eq. 9, acts as a regularizer to keep the local model gradients aligned with the global model gradients, thereby preserving model coherence despite non-IID data.
By dynamically adjusting the regularization parameter $\beta$ based on the performance difference between local and global models, the adaptive loss function enhances the alignment of local models with the global model, thereby improving generalization and performance in non-IID scenarios. When local performance is less compared to global, the regularization term $\beta$ amplifies the focus on local optimization (first term in Eq. 5), enabling better-performing clients to explore local optima more effectively. Conversely, if the global model performs better, this term shifts the emphasis towards aligning with the global model (second term in Eq. 5), thereby supporting clients that are struggling by incorporating global knowledge.
@@
[…]
@@
where $\bar{p}$ is the current estimate of the barycenter, $p_{i}$ refers to the $i^{th}$ client’s gradient distribution, $\epsilon$ is a small positive constant, and the iterations continue until convergence. After few iterations, we get the Wasserstein barycenter that is used to update the global model weights. We update the the global model weights for the last layers by substracting them from the calculated Wasserstein barycenter for effectively aggregating the updates from the last layers. The algorithm of proposed method FedDUAL is given in the Algorithm  1. The proof of the convergence for both convex and non-convex settings for the proposed method can be found in the Section  B of the Appendix.
@@
3:Define a mask $e\in\{0,1\}^{d}$, where $e_{j}=1$ for the last few layers and 00 for the rest layers.
@@
12:         $\mathcal{W}\leftarrow\mathcal{W}\cup\{\theta_{k}\}$triangleright Store local model weights $\theta_{k}$
@@
14:         $\Delta\leftarrow\Delta\cup\{\nabla_{k}\}$triangleright Store gradients $\nabla_{k}$
@@
19:              Compute Wasserstein Barycenter of last layer $j$ gradients ¯$\bar{\nabla}_{j}$
@@
27:$\mathcal{G}^{*}\leftarrow\theta^{g}$triangleright Final trained global model

### v2

@@ Abstract
Federated Learning (FL) marks a transformative approach to distributed model training by combining locally optimized models from various clients into a unified global model. While FL preserves data privacy by eliminating centralized storage, it encounters significant challenges such as performance degradation, slower convergence, and reduced robustness of the global model due to the heterogeneity in client data distributions. Among the various forms of data heterogeneity, label skew emerges as a particularly formidable and prevalent issue, especially in domains such as image classification. To address these challenges, we begin with comprehensive experiments to pinpoint the underlying issues in the FL training process, such as gradient instability and the emergence of sharp minima in the global model, both of which contribute to performance inconsistencies. Based on our findings, we introduce an innovative dual-strategy approach designed to effectively resolve these issues. First, we introduce an adaptive loss function for client-side training, meticulously crafted to preserve previously acquired knowledge while maintaining an optimal equilibrium between local optimization and global model coherence. Secondly, we develop a dynamic aggregation strategy for aggregating client models at the server. This approach adapts to each client’s unique learning patterns, effectively addressing the challenges of diverse data across the network. Our comprehensive evaluation, conducted across three diverse real-world datasets, coupled with theoretical convergence guarantees, demonstrates the superior efficacy of our method compared to several established state-of-the-art approaches. The code can be found at https://github.com/Pranabiitp/FedDUAL.
@@ 1 Introduction
Federated learning (FL) has revolutionized collaborative model training by enabling multiple clients to contribute to a global model without compromising the privacy of their local data (McMahan et al., 2017). This decentralized strategy avoids the need for sending data to a central server, thus maintaining data privacy. As the digital landscape evolves, with an increasing number of distributed data sources emerging from mobile devices, healthcare institutions, and Internet of Things (IoT) networks, FL has emerged as a pivotal solution for training sophisticated deep networks across geographically dispersed and heterogeneous environments (Bonawitz et al., 2016)Sahoo et al. (2024a)(Hu et al., 2024). However, a significant practical obstacle encountered during federated training is data heterogeneity in the form of skewness in labels and quantity of the data across various clients (Kairouz et al., 2021)(Li et al., 2020). Diverse user behaviors can lead to significant heterogeneity in the local data of different clients, leading to non-independent and identically distributed (non-IID) data. This variability can introduce biases in model training, leading to unstable convergence and potentially degrading the model’s performance or making it counterproductive (Li et al., 2022)(Zhao et al., 2018). While FedAvg (McMahan et al., 2017) is effective and widely used, it often falls short in accuracy and convergence with static aggregation methods. These methods combine model updates from different clients in a fixed manner, failing to adapt to heterogeneous data distributions and client drift, as discussed in (Karimireddy et al., 2020).
Previous studies have addressed the issue of client drift by implementing penalties for deviations between client and server models (Li et al., 2020)(Li et al., 2021a), employing variance reduction techniques during client updates (Karimireddy et al., 2020)(Acar et al., 2021), or utilizing novel aggregation methods on the server side (Chen et al., 2023)(Chowdhury and Halder, 2024).
@@ 1.1 Motivation
Prior studies by Yashwanth et al. (2024)Hu et al. (2024) have demonstrated that in non-IID scenarios, federated models tend to converge to ’sharp minima’, resulting in significant performance degradation and compromised generalizability. In this study, we investigate the root causes of this phenomenon and propose a novel solution to mitigate its effects. Our study begins with a detailed analysis of loss landscapes for FedAvg-trained models across IID and non-IID data distributions. Figure 1 visually depicts the loss landscapes of two models on the FMNIST dataset with systematic parameter perturbations. The model trained on IID data exhibits a notably smoother and wider valley in its loss landscape, suggesting greater robustness and better generalization. In contrast, the model trained on non-IID manifests sharper peaks and narrower valleys, indicating higher sensitivity to parameter variations and potential overfitting. These visualizations offer strong evidence that in the presence of non-IID data, the FedAvg algorithm achieves suboptimal generalization. Motivated by this observation, we investigate the underlying mechanisms by analyzing gradient norms to identify which parts of the neural network are most affected by data heterogeneity. Our findings, presented in Fig. 2 reveal a notable pattern: in non-IID scenarios, the gradient norms of the final layers, including the classification layer, exhibit significant amplification compared to their IID counterparts. Such amplification leads to model instability, impedes convergence, and ultimately compromises the generalizability of the federated model. Our investigation suggests that effective federated training in non-IID environments necessitates targeted adjustments during server-side aggregation, particularly for these highly affected layers, to achieve performance comparable to IID settings.
@@
This prompts one critical question: Can static aggregation methods effectively address severe non-IID data distributions across clients while maintaining higher convergence, performance, and generalizability in federated models? The answer is decidedly negative. Static aggregation methods inherently struggle with the dynamic heterogeneity present in federated networks, where adjusting parameters based on client distributions and performance in each communication round is crucial. Although incorporating predetermined parameters into the aggregation process may provide some partial mitigation, these methods fail to address the complex challenges posed by non-IID data distributions. A more dynamic and nuanced approach is necessary to effectively manage these multifaceted issues. To address this challenge, we apply dynamic aggregation to the model’s final layers, where gradient norms fluctuate significantly in non-IID scenarios, while using traditional aggregation (FedAvg) for the lower layers. For dynamic aggregation, we leverage the concept of Wasserstein Barycenter (Agueh and Carlier, 2011), derived from optimal transport theory, to integrate client-specific learning behaviors in these affected layers. By minimizing discrepancies from non-IID data, the Wasserstein Barycenter helps to align gradients from diverse clients, offering precise model updates. This approach ensures fair aggregation, adapts to data heterogeneity, reduces bias, and enhances robustness, ultimately leading to more stable model convergence and improved generalization.
@@
Figure 2: Comparison of gradient norms between models trained on IID and non-IID datasets using the FedAvg algorithm. (a) FMNIST dataset using LeNet model. (b) CIFAR10 dataset using VGG16 model.
@@ 2 Related Work
Another line of research targets optimizing server-side aggregation in FL. For instance, Hsu et al. (2019) investigated the impact of non-IID data on visual classification by creating datasets with diverse distributions and found that increased data heterogeneity negatively affected performance, leading them to propose server momentum as a potential solution. FedNova (Wang et al., 2020b) addressed the problem of objective inconsistency due to client heterogeneity in federated optimization by introducing a normalized averaging technique, which resolves this inconsistency and ensures rapid error convergence. Addressing the limitations of traditional parameter averaging methods, Lin et al. (2020) introduced ensemble distillation for model fusion. This approach allows for the flexible aggregation of heterogeneous client models by training a central classifier on unlabeled data, using the outputs from the client models as guidance. FedMRL (Sahoo et al., 2024b) introduced a novel framework by using a loss function that promotes fairness among clients and employed a multi-agent reinforcement learning for personalized proximal terms , and a self-organizing map to dynamically adjust server-side weights during aggregation.
3 Definitions
In this subsection, we summarize the key mathematical definitions used in the paper to ensure clarity.
Kullback–Leibler (KL) Divergence.
The KL divergence measures how one probability distribution diverges from a second reference distribution. For two distributions $P$ and $Q$ over the same probability space, it is defined as:
$D_{\mathrm{KL}}(P\,\|\,Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}.$ (1)
Wasserstein Distance.
The Wasserstein distance (also known as the Earth Mover’s Distance) measures the optimal cost of transporting mass to transform one probability distribution into another. For two distributions $\mu$ and $\nu$, the $p$-Wasserstein distance is defined as:
$W_{p}(\mu,\nu)=\left(\inf_{\gamma\in\Gamma(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{X}}d(x,y)^{p}\,d\gamma(x,y)\right)^{1/p},$ (2)
where $\Gamma(\mu,\nu)$ denotes the set of all couplings with marginals $\mu$ and $\nu$, and $d(\cdot,\cdot)$ is a ground metric. We leverage the Wasserstein distance and its barycenter to aggregate client models in a manner that accounts for data heterogeneity.
Wasserstein Barycenter.
Given distributions $\{\mu_{k}\}_{k=1}^{K}$ and weights $\{\lambda_{k}\}_{k=1}^{K}$, the Wasserstein barycenter $\hat{\mu}$ is defined as the distribution minimizing the weighted sum of Wasserstein distances:
$\hat{\mu}=\arg\min_{\nu}\sum_{k=1}^{K}\lambda_{k}W_{p}(\mu_{k},\nu).$ (3)
This barycenter allows us to aggregate the last-layer representations of client models more effectively than simple averaging, improving robustness to non-IID data.
@@ 4 Methods and Materials
$\theta^{*}=\operatorname*{arg\,min}_{\theta}F(\theta),\hskip 2.0ptF(\theta)=\frac{1}{K}\sum_{k}f_{k}(\theta),$ (4)
@@
where $p^{k}_{i}$ and $q_{i}$ are the probabilities associated with the $i^{th}$ component of the weight vectors. The local model must excel on local data while maintaining alignment with the global model to enhance overall generalization. This balance between minimizing local loss and aligning with the global model is defined as local adaptive function $\tilde{f}_{k}(\theta)$ in Eq. 8.
$\tilde{f}_{k}(\theta)=(1-\beta)*f_{k}(\theta)+\beta*D_{\text{KL}}(p^{k}\|q),$ (8)
where $f_{k}(\theta)$ is cross-entropy loss for $k^{th}$ client and $\beta$ is a regularization parameter and should be adaptive to account for the performance discrepancy between the local and global models. When the local model substantially outperforms the global model, $\beta$ should increase to enforce greater alignment. Conversely, if the models perform similarly, $\beta$ should decrease, allowing the local model to focus more on local optimization. The definition of $\beta$ is given in Eq. 9.
@@
$w_{k}^{t+1}=w_{k}^{t}-\eta\nabla_{w}\mathcal{L}_{\text{adaptive}}^{k}(w_{k}^{t}),$ (11)
where $\eta$ is the local learning rate. Expanding the gradient term $\nabla_{w}\mathcal{L}_{\text{adaptive}}^{k}(w_{k}^{t})$, we obtain Eq. 12:
$\nabla_{w}\mathcal{L}_{\text{adaptive}}^{k}(w_{k}^{t})=(1-(\sigma(\mathcal{A}_{\text{local}}^{k}-\mathcal{A}_{\text{global}}^{k}))\nabla_{w}\mathcal{L}_{\text{local}}^{k}(w_{k}^{t})+\sigma(\mathcal{A}_{\text{local}}^{k}-\mathcal{A}_{\text{global}})\nabla_{w}D_{\text{KL}}(p^{k}\|q).$ (12)
The KL divergence term, $\sigma(\mathcal{A}_{\text{local}}^{k}-\mathcal{A}_{\text{global}}^{k})\nabla_{w}D_{\text{KL}}(p^{k}\|q)$ in Eq. 12, acts as a regularizer to keep the local model gradients aligned with the global model gradients, thereby preserving model coherence despite non-IID data. The adaptive coefficient $\beta$ (Eq. 9) is dynamically computed as a function of the performance gap $(A^{k}_{\text{local}}-A^{k}_{\text{global}})$. When the global model outperforms the local one, $\beta$ tends toward 0 (via the sigmoid function), thus increasing the weight on the local loss term $(1-\beta)$ and enabling the client to focus more on its local data. Conversely, when the local model performs better than the global, $\beta$ increases toward 1, strengthening the KL term to preserve global knowledge rather than blindly aligning the models. The rationale is that superior local accuracy often reflects overfitting to non-IID client data. A higher $\beta$ in such cases regularizes the local model by constraining it to remain close to the global parameter manifold, improving overall generalization. Conversely, when the global model generalizes better, a smaller $\beta$ allows the client to emphasize local learning. For example, a client trained only on classes 0–4 may achieve high local accuracy but would poorly generalize to unseen classes 5–9. The higher $\beta$ and associated KL regularization preserve the global model’s multi-class knowledge, thereby preventing catastrophic forgetting.
Note that we employed KL divergence on the softmax of the parameters to capture distributional alignment between local and global models rather than direct numerical differences. By transforming flattened parameters into probability distributions through softmax, we interpret model weights as expressing relative importance rather than absolute magnitude. KL divergence thus quantifies how the structural configuration of model parameters diverges between clients and the server. Compared to alternatives, logit-based distances capture only output-level similarity and overlook the underlying parameter drift that degrades generalization. Layer-wise representation distances would require computing activations on a reference dataset, introducing computational overhead and ambiguity regarding data selection in heterogeneous settings. Similarly, the L2 parameter distance (as in FedProx) reflects magnitude differences but not distributional structure which is an essential factor for understanding model behavior. Empirically, we find KL divergence advantageous because its asymmetry aligns with the local to global adaptation objective and its bounded nature ensures stable gradients during optimization.
@@
$W(\mu_{k},\mu_{j})=\left(\inf_{\gamma\in\Gamma(\mu_{k},\mu_{j})}\int_{\mathcal{X}\times\mathcal{X}}d(x,y)^{p}\,d\gamma(x,y)\right)^{1/p}$ (14)
@@
where $\bar{p}$ is the current estimate of the barycenter, $p_{i}$ refers to the $i^{th}$ client’s gradient distribution, $\epsilon$ is a small positive constant, and the iterations continue until convergence. After few iterations, we get the Wasserstein barycenter that is used to update the global model weights. We update the the global model weights for the last layers by substracting them from the calculated Wasserstein barycenter for effectively aggregating the updates from the last layers. Our gradient analysis (Fig. 2) indicates that non-IID data disproportionately amplifies gradient norms in the final layers, leading to heightened instability compared to IID conditions. Traditional averaging methods aggregate client updates in Euclidean space without accounting for the underlying distributional differences across clients. The Wasserstein Barycenter addresses this limitation by operating in the space of probability distributions, finding the optimal aggregation point that minimizes distributional discrepancies (Wasserstein distances) to all client updates. This geometry-aware approach provides more robust aggregation by explicitly accounting for heterogeneous client learning behaviors, reducing the bias and instability caused by non-IID data. The algorithm of proposed method FedDUAL is given in the Algorithm  1. The proof of the convergence for both convex and non-convex settings for the proposed method can be found in Section A of the Appendix.
@@
3:Define a mask $e\in\{0,1\}^{d}$, where $e_{j}=1$ for the last few layers and 0 for the rest layers.
@@
12:   $\mathcal{W}\leftarrow\mathcal{W}\cup\{\theta_{k}\}$ $\triangleright$ Store local model weights $\theta_{k}$
@@
14:   $\Delta\leftarrow\Delta\cup\{\nabla_{k}\}$ $\triangleright$ Store gradients $\nabla_{k}$
@@
19:     Compute Wasserstein Barycenter of last layer $j$ gradients $\bar{\nabla}_{j}$
@@
[…]